OpenStudy (anonymous):

In a randomized trial, 200 subjects were to be randomly assigned either to the drug treatment group or to the placebo group. Researchers flipped a coin to decide each participant’s treatment assignment. After randomization, 112 subjects were in the treatment group and 88 were in the control group. What’s the probability that 112 or more subjects could end up in the treatment group purely by chance? Hint: use a normal approximation to the binomial.

OpenStudy (harsimran_hs4):

let`s start tell what is the probability of exactly 112 subjects getting assigned to treatment group ?

OpenStudy (anonymous):

50%

OpenStudy (harsimran_hs4):

tell me the method(explanation) you used

OpenStudy (anonymous):

flipping a coin.

OpenStudy (harsimran_hs4):

well you answer is not correct, let`s go step by step we have to decide for 1st subject, so we flip the coin what are the chances that this subject lands in the treatment side ?

OpenStudy (anonymous):

when we flip a coin 1/2 chances he would land in treatment side 1/2 chances he would land in other

OpenStudy (harsimran_hs4):

great now take just the second subject and we flip a coin to decide, what are the chances that this subject lands in the treatment side ?

OpenStudy (anonymous):

1/2

OpenStudy (harsimran_hs4):

great now if i have to consider the events simultaneously i.e i have 2 subjects and the destiny of the subjects will be decided by the flip of the coin what are the chances that both of these fall on the treatment side ?

OpenStudy (anonymous):

chances of 1 plus chances of 2

OpenStudy (harsimran_hs4):

no plus is used when they are complementary events like the occurrence of 1st subject on the treatment side rules out the occurrence of 2nd subject on the treatment side multiplication is used when both the sub events are part of the major event, or they are independent of each other like in our case 1st subject going into treatment side has not effect on the probability of the 2nd subject so now tell me the answer again?

OpenStudy (anonymous):

ok it would be 1/2 x 1/2 or .5 x .5

OpenStudy (harsimran_hs4):

yes now tell me what is the probability of getting exactly 112 subjects on treatment side?

OpenStudy (anonymous):

112 x .5?

OpenStudy (harsimran_hs4):

no (0.5)*(0.5)*(0.5) so on 112 times (THINK WHY??) i.e (1/2)^112 where * = multiply ^ = raised to power tell me when this step is clear as this is very important

OpenStudy (anonymous):

yes got that.clear

OpenStudy (harsimran_hs4):

ok now tell me the probability for exactly 113 such cases ?

OpenStudy (anonymous):

.5x.5x.5............113 times

OpenStudy (harsimran_hs4):

yes now you should realize that occurrence of exactly 112 and exactly 113 is complementary to each other clear ?

OpenStudy (anonymous):

yes

OpenStudy (harsimran_hs4):

so probability of 112 or more landing on treatment side is (1/2)^112 + (1/2)^113 + ... +(1/2)^200 clear ?

OpenStudy (anonymous):

yes.clear till here

OpenStudy (harsimran_hs4):

so all you need now is a way to find the sum it`s a GP which ratio 2

OpenStudy (anonymous):

how to make now these big calculations and find a whole number percentage

OpenStudy (harsimran_hs4):

check this out http://en.wikipedia.org/wiki/Geometric_progression

OpenStudy (anonymous):

confused :(

OpenStudy (harsimran_hs4):

S = a + ar +ar^2 + ar^3 + ...+ar^(n-1) S = a(1-r^n)/(1-r) a = (1/2)^112 and r = 1/2

OpenStudy (anonymous):

what is n?

OpenStudy (harsimran_hs4):

you can find out ar^(n-1) = last term = (1/2)^200 a and r = you know so find out n